# REMI: Reconstructing Episodic Memory During Intrinsic Path Planning

### v1

@@ Abstract
Grid cells in the medial entorhinal cortex (MEC) are believed to path integrate speed and direction signals to activate at triangular grids of locations in an environment, thus implementing a population code for position. In parallel, place cells in the hippocampus (HC) fire at spatially confined locations, with selectivity tuned not only to allocentric position but also to environmental contexts, such as sensory cues. Although grid and place cells both encode spatial information and support memory for multiple locations, why animals maintain two such representations remains unclear. Noting that place representations seem to have other functional roles in intrinsically motivated tasks such as recalling locations from sensory cues, we propose that animals maintain grid and place representations together to support planning. Specifically, we posit that place cells auto-associate not only sensory information relayed from the MEC but also grid cell patterns, enabling recall of goal location grid patterns from sensory and motivational cues, permitting subsequent planning with only grid representations. We extend a previous theoretical framework for grid-cell-based planning and show that local transition rules can generalize to long-distance path forecasting. We further show that a planning network can sequentially update grid cell states toward the goal. During this process, intermediate grid activity can trigger place cell pattern completion, reconstructing experiences along the planned path. We demonstrate all these effects using a single-layer RNN that simultaneously models the HC-MEC loop and the planning subnetwork. We show that such recurrent mechanisms for grid cell-based planning, with goal recall driven by the place system, make several characteristic, testable predictions.
1Department of Electrical and Systems Engineering,
University of Pennsylvania, Philadelphia, PA 19104, USA
2Tel Aviv Sourasky Medical Center, Tel Aviv 6423906, Israel
3Gray Faculty of Medical and Health Sciences, Tel Aviv University, Tel Aviv 6997801, Israel
4Rappaport Faculty of Medicine, Technion – Israel Institute of Technology, Haifa 31096, Israel
5Department of Physics, University of Pennsylvania, Philadelphia, PA 19104, USA
6Rudolf Peierls Centre for Theoretical Physics, University of Oxford, Oxford OX1 3PU, UK
7Santa Fe Institute, Santa Fe, NM 87501, USA
@@ 1 Introduction
Mammals employ various cell types to represent space and guide navigation [1, 2, 3]. For example, grid cells (GCs) in the Medial Entorhinal Cortex (MEC) fire in hexagonally arranged patterns as an animal moves through an environment [4, 5, 6]. Grid cells also exhibit multiple spatial scales in their periodic firing patterns [7, 8, 9, 10], which may arise through self-organization driven by inhibitory gradients in MEC attractor networks [11], enabling the efficient encoding of a large number of locations [12]. Meanwhile, hippocampal place cells (HPCs) fire at selective locations within an environment in a seemingly random manner [1, 2]. However, they are tuned not only to spatial locations but also to experiences and contextual cues [13, 14, 15, 16, 17]. The HPC population responses remap to form distinct representations in different environments and reinstate prior maps in familiar contexts. This ability supports their capacity to encode multiple environments without interference [13, 2, 18, 19, 20].
The local circuitry driving single GC and HPC activity has been studied extensively, and recent work shows that many of the phenomena they manifest can emerge in trained neural networks modeling their functional roles. Path integration theories propose that GCs track an animal’s position by integrating its movement, forming a continuous attractor landscape [4, 21, 22]. Likewise, recurrent neural networks (RNNs) trained to infer position from movement sequences develop grid-like patterns in their hidden layers [23, 24, 25, 26, 27]. For place cells, one view is that they could be emergent patterns that arise from encoding sequences of sensory signals relayed partly through the MEC [28, 29, 30] during spatial traversal [15, 16, 17, 31], perhaps implemented by a continuous attractor network (CAN) [32, 33, 34, 35]. An alternative account is the successor representation framework [36, 37], which suggests that HPCs may encode predictive representations of future spatial states [38, 39].
Building on extensive studies of the emergence and functional roles of grid and place cells [40, 23, 25, 27, 15, 16, 17, 41], a natural view is that animals maintain these representations not just to encode space, but to support intrinsically motivated navigation tasks such as recalling goal locations from partial cues and planning paths to them. This view is supported by evidence that during rest or sleep, the hippocampus reactivates neural sequences that either mirror past paths (replay) or form novel trajectories through unvisited locations (preplay) [42, 43]. Yet, theories of place cell emergence suggest their dynamics are updated by sensory input sequences. Any planning network driving replay or preplay must therefore reproduce these inputs, or at least the sequence of sensory latent representations along the trajectory. However, if such a network could directly store and recall detailed sensory information along trajectories, the role of place cells in recall would be redundant, contradicting their well-established importance in spatial memory
One candidate framework suggests that predictive information encoded by place cells supports transitions from the current location to neighboring states. This encoded transition likelihood traces out a most likely path connecting the current and goal locations [37]. However, this framework requires the animal to visit and encode all locations in the environment, and therefore cannot explain how animals might take shortcuts through unvisited locations. An alternative idea leverages residue number systems, enabled by multiple scales of grid cells, to encode pairwise displacements [44]. However, while grid cells can represent a vast number of locations, encoding pairwise relations among them quickly becomes intractable, fails to account for hippocampal replay and preplay, and does not explain how grid patterns would be recalled to support navigation.
Inspired by recent work suggesting that place cells may autoassociatively encode sensory inputs [16], or more generally, signals weakly modulated by space, we propose a framework unifying grid-cell-based planning with contextual recall in the hippocampus. The spatial stability of grid representations [45] allows us to treat grid cell activity as a special form of observation derived from self-motion and modulated by space. Given that the MEC, which projects to the hippocampus, contains both grid cells and neurons relaying diverse sensory signals [28, 29, 30], we propose that if place cells autoassociate both, then: (1) partial sensory input will reactivate the corresponding grid cell state through hippocampal autoassociation, enabled by recurrent connections between HC and MEC; (2) a planning network operating in grid space can use stable relative phase relationships across rooms to plan paths without rewiring when entering new environments; (3) the known local transition rules of grid cells will enable the network to construct valid paths even through unvisited areas; and (4) during planning, grid representations will advance to intermediate states, and the hippocampus will retrieve the associated sensory cues, enabling reconstruction of sensory experience along the planned trajectory.
@@ 2 Method
2.1 Hippocampus–Medial Entorhinal Cortex Loop
During spatial traversal, animals form internal representations of the world shaped by olfactory, visual, and auditory cues. These sensory signals are partially reflected by (weakly) spatially modulated cells (SMCs) in the MEC [28] and relayed to the hippocampus (HC) [46, 47]. Such observations signal the presence of, e.g., food, water, or hazards, and thus may serve not only as inputs but also as objectives for navigation. Animals also receive displacement signals from the vestibular system, proprioception, and visual flow. Although these signals may not directly correspond to specific objectives, they help animals infer their relative position from previous locations. This process is supported by grid cells (GCs) in the MEC through path integration [4, 48, 49, 22].
However, both sources of information can fail during navigation. Sensory observations may be unavailable in conditions such as dim light or nighttime, while path integration may break down due to loss or corruption of speed or direction signals. We hypothesize that animals can exploit spatial correlations between these two sources of information to achieve more accurate localization than either source alone. However, given that these correlations are context-dependent, for example varying across compartments of an enclosure or configurations of natural landmarks, this relationship is unlikely to be supported by fixed wiring between SMCs and GCs within the MEC. Instead, we propose that hippocampal place cells (HPCs) encode these correlations through pattern completion, allowing the two types of information to be flexibly wired across contexts.
With this coupling between SMCs and GCs through HPCs, animals can directly retrieve a location’s GC representation from sensory cues. The animal can then use this recalled GC representation to plan a path, as planning with GC representations is not context dependent and can therefore be generalized to any environment [44]. Along a GC-planned path, this coupling could, in turn, reconstruct SMC activity from intermediate GC states. Since GCs maintain stable relative phase relationships across environments, we further propose that planning based solely on GC representations enables shortcuts through unvisited locations and immediate planning when re-entering familiar rooms. These capabilities are not supported by HPCs alone, as they encode discrete locations and lack the continuous spatial structure required for planning. To test this proposal, we construct an RNN model of the HC-MEC loop and build a planning model on top of it.
2.1.1 A RNN Model of the HC-MEC Loop
@@
Figure 1: (a) RNN model of the HC-MEC loop. The top subnetwork contains HPCs, with example emergent place fields shown on the left. The bottom subnetwork includes partially supervised GCs, as well as supervised speed, direction, and spatially modulated cells (SMCs); example grid fields shown at left. (b) & (d) Speed cells and direction cells are denoted as Spd and Dir, respectively. Colored regions highlight within-group recurrent connectivity to indicate the partitioning of the connectivity matrix by cell groups. However, at initialization, no structural constraints are enforced. The full connectivity matrix is randomly initialized. (b) Illustration of the path-integration network’s connectivity matrix. (c) The network is trained to path-integrate 5s trials and tested on 10s trials ($L=100.50\pm 8.49$ cm); the grid fields remain stable even in trials up to 120s ($L=1207.89\pm 30.98$ cm). For each subpanel (10s, 120s): top row shows firing fields; bottom row shows corresponding autocorrelograms. (d) Illustration of RNN connectivity matrix of full HC-MEC loop. (e-f) Example place fields (emergent) and grid fields in the full HC-MEC RNN model.
Integrate input and output cells directly into recurrent dynamics. Previous RNN models characterizing emergence of GCs and HPCs [23, 25, 27, 16, 50] have a limitation for planning: they typically do not explicitly model signals driving recurrent dynamics but instead rely on a learnable projection matrix. We posit that, during planning, these signals serve as control inputs to drive the HC-MEC loop toward the goal (see Sec. 4). Explicitly modeling these inputs simplifies the architecture and enables joint modeling of HC, MEC, and the planning subnetwork within a single recurrent structure.
Consider a standard RNN used in previous studies [23, 25, 27, 50, 16], updating its dynamics as:
$z_{t+1}=\alpha\cdot z_{t}+(\mathbf{1}-\alpha)\cdot\left(\mathbf{W}^{in}u_{t}+\mathbf{W}^{rec}f(z_{t-1})\right)$ (1)
@@
$\begin{bmatrix}z_{t+1}\\ z^{I}_{t+1}\\ z^{O}_{t+1}\end{bmatrix}=\alpha\odot\begin{bmatrix}z_{t}\\ z^{I}_{t}\\ z^{O}_{t}\end{bmatrix}+(\mathbf{1}-\alpha)\odot\left(\begin{bmatrix}\mathbf{0}\\ u_{t}\\ \mathbf{0}\end{bmatrix}+\begin{bmatrix}\mathbf{W}^{rec}&\tilde{\mathbf{W}}^{in}&\mathbf{W}^{(13)}\\ \mathbf{W}^{(21)}&\mathbf{W}^{(22)}&\mathbf{W}^{(23)}\\ \tilde{\mathbf{W}}^{out}&\mathbf{W}^{(32)}&\mathbf{W}^{(33)}\end{bmatrix}f\left(\begin{bmatrix}z_{t}\\ z^{I}_{t}\\ z^{O}_{t}\end{bmatrix}\right)\right)$ (2)
@@
We assign speed cells and allocentric direction cells as input nodes that only receive inputs. The SMCs are set to both input and output nodes, trained to match the simulated ground truth. These SMCs are assumed to primarily respond to sensory cues during physical traversal. Supervision constrains their dynamics to reflect tuning to these signals, while learned recurrent connections formed during training are intended to reflect Hebb-like updates in brain that preserve this tuning structure. The ground truth signals for all cells are strictly positive to reflect firing rates (see Suppl. 3 for how ground truth signals are simulated). For direction cells, we assign allocentric preferred directions uniformly over $[0,2\pi)$ with a fixed angular tuning width, ensuring their responses remain non-negative and their population activity lies on a 1D ring.
Partially Supervise Grid Cells. For our proposed planning mechanism, we must model a complete HC-MEC loop containing stable patterns of both GCs and HPCs. However, no existing model exhibits the simultaneous emergence of grid and place cells. To address this, we supervise GCs to learn path integration. Specifically, we simulate GC population responses within a room and use them as ground truth. At the start of each training trajectory, RNN hidden units modeling grid cells are initialized to the ground-truth GC responses at the starting location. Along this trajectory, the network receives only speed and direction signals, directly input to the corresponding cells and relayed to the GC subpopulation through recurrent projections. We collect GC subpopulation states over time and penalize their deviation from simulated ground-truth responses, encouraging controllable GC activity while learning path integration. We refer to this as partial supervision and apply it to GCs for three reasons: (1) our primary focus is to test planning by the HC-MEC loop rather than the co-emergence of both GCs and HPCs; (2) HPC remapping in novel environments involves extensive reorganization of spatial tuning and synaptic connections driven by sensory input [51, 52, 40], making it harder to parameterize and simulate than GCs, which exhibit greater stability [45] and maintain stable relative phase relationships across environments [3, 53]; (3) the planning mechanism critically depends on the experimentally established stability of GC phase relationships and multiple spatial scales, as we will show in the following sections; these properties are difficult to control in existing emergence models.
We recognize that, biologically, HPCs appear before grid cells (GCs) [54, 55, 56, 3]; however, HPC firing fields become more spatially precise as GCs mature, suggesting an iterative refinement process [57, 56]. Conceptually, our HC-MEC model with partially supervised GCs can be seen as capturing this refinement phase, where the emergence of GCs enhances the spatial specificity of HPCs. Reflecting this process, we observe a quick emergence during early training and a corresponding reduction in HPC firing field widths after GCs are learned (see Suppl. 5.1).
@@
Testing GC Path Integration and HPC Emergence. Planning requires stable GC and/or PC representations. Before discussing planning, we first train the network to develop these representations. We first test whether a GC network trained with partial supervision can perform path integration. We simulate six modules with spatial periodicities scaled by the theoretical optimal factor $\sqrt{e}$ [12]. The smallest grid spacing is set to 30 cm, defined as the distance between two firing centers of a grid cell [58]. The grid spacing to field size ratio is 3.26 [59], with firing fields modeled as Gaussian blobs with radii equal to two standard deviations. We train this model on short random trajectories (5 s) but accurately path-integrate over significantly longer trajectories (10 s, 120 s) during testing (Figure 1c).
We then test the full HC-MEC loop model. Both supervised and partially supervised cells receive masked input and initial states, with supervised cells additionally receiving inputs at all timesteps. The HPC subnetwork receives signals only from MEC through recurrent connections and does not take external input. All cells are modeled within a single recurrent network without explicit connectivity constraints (see Figure 1d for initialization details). We observe HPCs emerge in the network, while GCs learn to path integrate (Figure 1e-f).

### v2

# REMI: Reconstructing Episodic Memory During Internally Driven Path Planning

@@ Abstract
Grid cells in the medial entorhinal cortex (MEC) and place cells in the hippocampus (HC) both form spatial representations. Grid cells fire in triangular grid patterns, while place cells fire at specific locations and respond to contextual cues. How do these interacting systems support not only spatial encoding but also internally driven path planning, such as navigating to locations recalled from cues? Here, we propose a system-level theory of MEC-HC wiring that explains how grid and place cell patterns could be connected to enable cue-triggered goal retrieval, path planning, and reconstruction of sensory experience along planned routes. We suggest that place cells autoassociate sensory inputs with grid cell patterns, allowing sensory cues to trigger recall of goal-location grid patterns. We show analytically that grid-based planning permits shortcuts through unvisited locations and generalizes local transitions to long-range paths. During planning, intermediate grid states trigger place cell pattern completion, reconstructing sensory experiences along the route. Using a single-layer RNN modeling the HC-MEC loop with a planning subnetwork, we demonstrate these effects in both biologically grounded navigation simulations using RatatouGym and visually realistic navigation tasks using Habitat Sim. Codes for experiments, simulations, and vision encoder are available at  [^1] Project Page: https://zhaozewang.github.io/remi , [^2] RatatouGym (Simulation Suite): https://ratatougym.github.io, [^3] Bottleneck MAE: https://github.com/grasp-lyrl/btnk_mae.
1Dept. of Electrical and Systems Eng., Univ. of Pennsylvania
2Tel Aviv Sourasky Medical Center
3Gray Faculty of Medical and Health Sciences, Tel Aviv University
4Rappaport Faculty of Medicine, Technion – Israel Institute of Technology
5Dept. of Physics, Univ. of Pennsylvania     6Santa Fe Institute
7Rudolf Peierls Centre for Theoretical Physics, University of Oxford
@@ 1 Introduction
Grid and place cells in the medial entorhinal cortex (MEC) and hippocampus (HC) form complementary spatial codes for navigation [1, 2, 3]. Grid cells (GCs) in MEC fire in periodic triangular patterns as animals move through space [4, 5, 6], spanning multiple spatial scales [7, 8, 9, 10] potentially shaped by inhibitory gradients within attractor networks [11]. Path integration theories propose that GCs track position by integrating movement [4, 12, 13], and recurrent neural networks (RNNs) trained for this task similarly develop grid-like activity [14, 15, 16, 17, 18]. Hippocampal place cells (HPCs) fire at specific locations [1, 2] and remap across contexts [19, 20, 21, 22, 23], forming sparse, orthogonal representations that minimize contextual interference [24, 25, 26]. They auto-associate sensory sequences relayed through MEC during exploration [27, 28, 29, 30], linking related memories via continuous attractor dynamics [31, 32, 33, 34], and emerge in RNNs that auto-encode spatial experience [21, 22, 23].
Despite extensive study of grid and place cell function [35, 14, 16, 18, 21, 22, 23, 36], it remains unclear how these representations support internally driven navigation, such as recalling goals from partial cues and planning novel routes. This ability is fundamental to flexible, goal-directed behavior, allowing animals to adapt to novel situations without extensive relearning. Evidence from rodent rest and sleep shows that hippocampal sequences replay past trajectories and generate novel ones through unvisited locations [37, 38], suggesting an underlying mechanism for offline planning.
Two complementary frameworks have been proposed to explain planning. The successor representation (SR) posits that hippocampal place cells encode expected future occupancy [39, 40], enabling flexible replanning when goals change [41, 42] and supporting generalization across related tasks through reusable predictive representations [43]. However, SR models discretize space into learned transition states, requiring experience at many locations and limiting interpolation to unvisited areas. Grid-cell models instead propose continuous, periodic codes that support vector-based navigation [44], with stable phase relationships providing context-independent spatial metrics. Yet encoding pairwise relations among locations becomes intractable as environments scale, and pure grid-based planning cannot easily capture context-specific constraints or hippocampal replay phenomena. Each framework offers a partial explanation of the problem, but how the brain integrates context-dependent memory with context-independent spatial metrics remains unresolved.
Here, we propose a unified theory of hippocampal-entorhinal circuitry that explains how internally driven navigation arises from the interaction between place and grid cells. This specific hippocampal-entorhinal wiring can enable cue-triggered goal retrieval and grid-based planning. Building on recent work showing that place cells autoassociatively encode sensory inputs [22], we propose that they also associate these inputs with GC activity. This coupling supports bidirectional pattern completion, whereby sensory cues can reactivate corresponding grid representations to guide planning, and grid states can reconstruct the expected sensory experiences along planned routes.
We verified these effects using a single RNN trained to autoencode sensory inputs, integrate velocity signals for path integration, and form autoassociative links between sensory and grid representations. Through training, the RNN developed an internal representation of the navigable space. To enable planning, we expanded the RNN’s hidden layer to include an additional planner subnetwork that drives the encoder’s dynamics with internally generated action sequences, enabling the system to traverse imagined paths between the current location and a recalled goal. To test whether this framework could generalize to realistic navigation, we replaced simulated sensory patterns with visual features from panoramic images in a photorealistic environment (Habitat-Sim [45, 46, 47]), encoded using a modified Masked Autoencoder (MAE) [48]. During planning, as the network traversed imagined trajectories, intermediate sensory states were decoded by the MAE into images that closely matched the expected views at corresponding locations.
Our theory predicts that (a) during planning, HC-MEC coupling will induce “look forward” sweeps in MEC spatial representation, resembling recent experimental results [49], (b) when presented with sensory cues associated with distant locations, grid cell response patterns corresponding to those locations will be reactivated, and (c) disrupting MEC-to-HC projections should impair goal-directed navigation, while disrupting HC-to-MEC feedback should reduce planning accuracy.
@@ 2 Method
Conceptual Framework of the HC-MEC Loop. We construct a conceptual framework showing how known spatial cell types could integrate into a unified system supporting both mapping and internally driven planning. Our theory builds on four established theories: (a) GCs arise from integrating velocity sequences for localization [14, 15, 16, 17, 18, 4, 12, 13]; (b) spatially modulated cells (SMCs) in MEC respond to external sensory stimuli [27] and relay sensory inputs to hippocampus; (c) HPCs autoassociate noisy sensory inputs from SMCs to reconstruct denoised sensory experience during recall [50, 51, 52, 21, 22]; and (d) while HPCs may also support planning, GCs provide a more efficient substrate for planning through their continuous, context-independent encoding [44].
@@
Figure 1: (a) RNN model of the HC-MEC loop. The top subnetwork contains HPCs, with example emergent place fields shown on the left. The bottom subnetwork includes partially supervised GCs, as well as supervised speed, head direction, and spatially modulated cells (SMCs); example grid fields shown at left. (b) & (d) Speed cells and head direction cells are denoted as Spd and Dir, respectively. Colored regions highlight within-group recurrent connectivity to indicate the partitioning of the connectivity matrix by cell groups. However, at initialization, no structural constraints are enforced. The full connectivity matrix is randomly initialized. (b) Illustration of the path-integration network’s connectivity matrix. (c) The network is trained to path-integrate 5s trials and tested on 10s trials ($L=100.50\pm 8.49$ cm); the grid fields remain stable even in trials up to 120s ($L=1207.89\pm 30.98$ cm). For each subpanel (10s, 120s): top row shows firing fields; bottom row shows corresponding autocorrelograms. (d) Illustration of RNN connectivity matrix of full HC-MEC loop. (e-f) Example place fields (emergent) and grid fields in the full HC-MEC RNN model.
Since both GCs and SMCs are present in MEC and both project to hippocampus, GCs may be regarded as a special form of sensory response derived from proprioceptive signals. This anatomical arrangement suggests our core theoretical contribution, extending (c) by proposing HPCs autoassociate and pattern-complete not only SMCs but also GCs. When sensory cues are unreliable and path integration drifts over time due to noise, this association provides a natural solution for more robust localization than either system alone. This framework directly yields three key predictions: (1) partial sensory cues can reactivate the corresponding grid cell state via hippocampal autoassociation; (2) the recalled grid state can guide planning on the grid cell manifold, enabling efficient, context-independent, and generalizable path computation; (3) during planning, intermediate grid states trigger hippocampal reactivation of the corresponding sensory representations, reconstructing the expected sensory experience along the planned route.
An RNN Model of the HC-MEC Loop. Our theoretical framework requires an RNN that can simultaneously (1) autoencode masked and noisy sensory experiences (SMCs), (2) contain grid cells capable of path-integrating noisy movement inputs, (3) form autoassociative links between SMC and GC responses, and (4) later support planning. To achieve these, we extend RNN formulations from previous models of spatial navigation cell types [14, 16, 18, 22, 53]. Specifically, consider a standard RNN that updates its dynamics as:
$z_{t+1}=\alpha\cdot z_{t}+(\mathbf{1}-\alpha)\cdot\left(\mathbf{W}^{in}u_{t}+\mathbf{W}^{rec}f(z_{t})\right)$ (1)
@@
[…]
@@
Supervising Spatially Modulated Cells (SMCs). The SMCs are set as both input and output nodes, trained to reproduce the simulated ground truth. At each time step, SMC units receive simulated but partially occluded sensory inputs representing noisy environmental observations. After each recurrent update, the network is expected to reconstruct the corresponding noiseless sensory pattern and is penalized for deviations of the SMC subpopulation from the ground truth responses. These SMCs are assumed to primarily respond to sensory cues during physical traversal. Supervision constrains their dynamics to reflect tuning to these signals, while learned recurrent connections formed during training are intended to reflect Hebb-like updates in the brain that preserve this tuning structure.
Supervising Grid Cells for Path-Integration. Our theoretical model also requires grid cells that have learned to perform path integration. To achieve this, we assign two additional subpopulations within the RNN’s hidden units to represent speed and allocentric head direction cells. Both subpopulations are defined as input nodes that only receive external inputs. Head direction cells are assigned preferred allocentric directions uniformly distributed over $0,2\pi)$ with a fixed angular tuning width, ensuring non-negative responses and a population activity that lies on a 1D ring (see Suppl. 3). We also simulate ground-truth grid cell activity maps. To train the network to perform path integration, the model is provided only with the initial grid cell responses at $t=0$. For all subsequent timesteps, the network must infer the grid cell activity from the initial response together with the ongoing inputs from the speed and head direction cells.
@@
This path integration component can be viewed as a modified version of the networks described in [14, 16, 15, 18]. To verify that the network indeed learned path integration, we first trained a model containing only grid, speed, and head direction cells (Figure 1b). We simulated six grid modules with spatial periods scaled by the theoretical optimal factor $\sqrt{e}$ [54]. The smallest grid spacing is set to 30 cm, defined as the distance between two firing centers of a grid cell [55]. The grid spacing to field size ratio is 3.26 [56], with firing fields modeled as Gaussian blobs with radii equal to two standard deviations. We train this model on short random trajectories (5 s) but accurately path-integrate over significantly longer trajectories (10 s, 120 s) during testing (Figure 1c).
Emergence of Place Cell-Like Patterns During Autoassociation of SMCs and GCs. The central idea of our HC-MEC model is that localization using sensory inputs (SMCs) or through path integration alone can each fail under noise. Associating the two representations through hippocampal place cells makes localization more robust. To model this, we introduced a subpopulation of hidden units in the RNN that are neither input nor output neurons. They do not directly receive external signals or produce outputs. These units receive input from and project to both the SMC and GC subpopulations through recurrent connections (Figure 1d). After training the full HC-MEC network, with SMC units supervised to autoencode masked sensory inputs and GC units trained to path integrate noisy movement signals, we observed the spontaneous emergence of place cell-like activity in this intermediary region. No explicit supervision or spatial constraint was imposed, and the place-like patterns emerged naturally through training, consistent with previous findings [22].